OpenStudy (anonymous):

implicit differentiation.. x^2 - 2xy + y^2 - 6x + 2y=0

OpenStudy (anonymous):

the answer is dy/dx = x-y-3/x-y-1.. it was written on the book

OpenStudy (anonymous):

\[ -2 x y'(x)+2 y(x) y'(x)+2 y'(x)-2 y(x)+2 x-6=0 \] Solve for y'(x) and you are done

OpenStudy (anonymous):

\[ -2 \left(x y'(x)-y(x) y'(x)-y'(x)+y(x)-x+3\right)=0\\ x y'(x)-y(x) y'(x)-y'(x)+y(x)-x+3=0\\ x y'(x)-y(x) y'(x)-y'(x)=-y(x)+x-3\\ y'(x)( x -y(x) -1)=-y(x)+x-3\\ y'(x)=\frac{-y(x)+x-3}{-y(x)+x-1} \]

OpenStudy (anonymous):

yay thank you so much @eliassaab

OpenStudy (anonymous):

YW @silverxx